the number of diagonals d of a polygon of n sides is given by d = n/2 (n-3). How many sides does a polygon with 77 diagonals have?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(53765)
Answer by Alan3354(69443)

the number of diagonals d of a polygon of n sides is given by d = n/2 (n-3). How many sides does a polygon with 77 diagonals have?
-------------------
(n/2)*(n-3) = 77
n^2 - 3n = 154
n^2 - 3n - 154 = 0
(n-14)*(n+11) = 0
n = 14 sides (Ignore the -11)